Sol Summary – December 8th

Crew 200 Sol Summary Report 08-DEC-2018

Sol: 6

Summary Title: First solar observation, mission extended for reduced
crew members

Author’s name: Lindsay Rutter

Mission Status: Mission is completed. A proposal was made to extend
the mission with a reduced crew number of three people who will work
toward a relatively new set of objectives. Proposal was approved.

Sol Activity Summary: We woke up today to the smell of lobster
biscuits and herb garlic butter. Several crew members had been
salivating over that box of lobster biscuits since Sol 1, so it was a
celebratory meal that marked the end of our nominal mission. The skies
were clear today and our crew astronomer was kind enough to show
several crew members how to use the solar observatory. After cleaning
and bag packing, we drove to Hanksville to share lunch together and
part ways. Four crew members continued to Grand Junction, while three
crew members returned to the habitat. It was hard to see the crew
split into half, but the three remaining crew members are determined
to use what they have learned from the full team to make the extended
mission worthwhile at both a personal level and team level.

Look Ahead Plan: The reduced crew members will plan new and modified
research projects and brainstorm possible reduced crew EVAs.

Anomalies in work: N/A

Weather: Warmer than recently, snow melting, clear and sunny skies.

Crew Physical Status: No crew members reported physical problems today.

EVA: N/A

Reports to be file: N/A

Support Requested: N/A

Operations Report – December 8th

Crew 200 Operations Report 08-Dec-2018

SOL: 7

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart, generator start, solar panels power generation or conversion
Notes on non-nominal systems:
1) During the day the solar panels almost generated no power (<1kW for 2, 0kW for the 3rd) although it was a very sunny day. This led to a drop of SOC to 66% when we came back from town (at 17:20).
2) At 17:20, we started the generator which worked on the generator side, however after pushing the button to connect the generator to the network, the system shuts down after a few minutes (it was in under frequency). We tried several times and it did the same. Shannon came to help and after repeating the operation several times, it finished to work.

Generator (hours run): From 4:00 pm last night to 7:45 am. On tonight at 6:15 pm.

Solar— SOC 65%
Diesel Reading – 25 %
Propane Reading – 38 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 300 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – not noted gallons, low.
Water (loft) – Static to Loft Pump used – yes
Water Meter: 6

Toilet tank emptied: yes

Deimos rover used: no, still not functional
Hours:
Beginning charge:
Ending charge: 73%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were not used this week
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? Yes, to go to Hanksville at 1:00 pm as a part of the crew is leaving today, we ate there and came back at 5:10
CrewCar used and why, where? Yes, by Shannon, to go to Hanksville at 1:00 pm as a part of the crew is leaving today, we ate there and she came back at 5:50

General notes and comments: the nominal mission is completed (1 week sim), we stopped the sim at 0:00am and cleaned up the hab. While 4 members of the crew left (with their own car), an extension of the mission was proposed and approved by Shannon, with a 3 member crew. We will start again the sim tomorrow evening.
Generator start was problematic.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Cleaning before half of the crew left

Summary of GreenHab operations: Regular activities, temperature was rised to 50 F.

Summary of ScienceDome operations: ScienceDome was not sufficiently cleaned, we will take care of this week

Summary of RAM operations: it is not leaking anymore from the roof (the snow melted)

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Greenhab Report – December 8th

Crew 200 Greenhab Report 08-DEC-2018

Greenhab Officer:
Makiah Eustice

Environmental Control:
Heating
40% Shade Cloth on

Average Temperatures:
Low: 19 C
High: 52 C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
9 gallons

Water in Blue Tank (lbs):
27%

Times of Watering for Crops:
1100, 1830

Changes to crops:
NONE

Narrative:
I realized that I am really underwatering my plants. The new crew of 3
spent time watering tonight. We thinned the carrots as well. Shannon
also increased the average temperature. We thinned the carrots as
well.We will be refilling the water tomorrow.

Harvest:
NONE

Support/Supplies Needed:
NONE